as x approaches infinity
find the limit of
e^ (x- x^2)

Answer by stanbon(75887)

as x approaches infinity
find the limit of
e^ (x- x^2)
----
e^(x-x^2) = e^x/e^x^2 = 1/e^x
--
lim as x approaches oo = 1/oo = 0
========================================
Cheers,
stan H.